OpenStudy (anonymous):

WILL GIVE MEDAL HELP ME please Find the measures of the angles. m∠ACD m∠CED A. m∠ACD = 40° m∠CED = 35° B. m∠ACD = 50° m∠CED = 45° C. m∠ACD = 50° m∠CED = 35° D. m∠ACD = 40° m∠CED = 45°

OpenStudy (anonymous):

m<ACD = 90 - 50 = 40 deg m<CED = 180 - 145 = 35 deg

OpenStudy (anonymous):

thank you SO MUCH